OpenStudy (anonymous):

If f(x) = 1/2 x + 2 and g(x) = −3x − 1, find g[f(4)]. −19 11 −13 −11

OpenStudy (kinggeorge):

If you're given something like g[f(x)], that means that you want to find f(x), and take that value, and plug into to g. For example: g[f(0)]: Since \(\displaystyle f(0)=\frac{1}{2}(0)+2=2 \) \( g[f(0)]=g(2)=-3(2)-1=-7\). Now can you do this with 4 instead of 0?

OpenStudy (anonymous):

so iht would be -11

OpenStudy (kinggeorge):

You're very close. You have f(4)=4 correct?

OpenStudy (anonymous):

i got it -13

OpenStudy (kinggeorge):

Bingo.

OpenStudy (anonymous):

thanx

OpenStudy (kinggeorge):

you're welcome.

OpenStudy (anonymous):

can yhu help with one other question to?